A Jet lands on a aircraft carrier at 180 mile (123 m/s) (a)what is its acceleration if it stops in 3 second (b)what is the displacement of the plane which it is stopping
1
Expert's answer
2020-02-17T09:15:30-0500

(a) We can find the acceleration of the plane from the kinematic equation:


"v_f = v_i + at,"

here, "v_i = 123 \dfrac{m}{s}" is the initial velocity of the plane, "v_f = 0 \dfrac{m}{s}" is the final velocity of the plane when it stops, "a" is the acceleration of the plane and "t" is the time interval during which the plane is stopping.

Then, from this equation we can find the acceleration of the plane:


"a = \dfrac{v_f-v_i}{t} = \dfrac{0 \dfrac{m}{s}-123 \dfrac{m}{s}}{3s} = -41 \dfrac{m}{s^2}."

The sign minus indicates that the jet is decelerates.

(b) We can find the displacement of the plane from another kinematic equation:


"d = \dfrac{1}{2}(v_i + v_f)t = \dfrac{1}{2} \cdot 123 \dfrac{m}{s} \cdot 3s = 184.5 m."

Answer:

(a) "a = -41 \dfrac{m}{s^2}."

(b) "d = 184.5 m."